OpenStudy (anonymous):

I am stuck on this question: A beta particle (fast moving electron) is released by a nuclear decay and starts travelling at 1.5x10^6 m/s [NΘE] in the Earth's magnetic field which has a value of 3.0x10^-5 T. If the force experienced by the beta particle is 4.34x10^-18 N, determine the angle between the direction of travel and the Earth's magnetic field. What is the direction of the force acting on the particle?

OpenStudy (anonymous):

Since \(\vec F = q(\vec v \times \vec B)\) so, \(\vec F = -evB sin \theta \hat n \) where \(\theta\) is angle between the direction of travel and the Earth's magnetic and \(\hat n\) is a unit vector perpendicular to the plane of \(\vec v \ and \ \vec B\). Put the given values and solve for \(\theta\).

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

is there a set value for e/q? (the value measured in Coulombs)

OpenStudy (anonymous):

I don't understand your question.. :(

OpenStudy (anonymous):

Oh sorry. In the formula f=qvBsin(theta) the f value would be 4.34x10^-18 N the B value would be 3.0x10^10^-5 T the v value would be 1.5x10^6 m/s the question asks for thr angle (sin theta) but I don't know what to put in place of q

OpenStudy (anonymous):

q is the charge of beta particle which is e = 1.6 * 10^-19